- 1. O, effuses twice as fast as an unknown gas X at the same temperature. What is the molar mass of the unknown gas? A. 8.0 the answer is E, but I don’t understand how. When I worked it out I got 8.0 g/mol
Answer #1

Rate of effusion is inversely proportional to square root of molar mass

rate = k*sqrt(1/M)

M is molar mass

If we take ratio, above expression becomes

rate(O2)/rate(unknown) = sqrt (M(unknown)/M(O2))

we have:

M(O2) = 32.0 g/mol

rate(O2) = 2.0

rate(unknown) = 1.0

rate(O2)/rate(unknown) = sqrt (M(unknown)/M(O2))

2.0/1.0 = sqrt (M(unknown)/32.0)

sqrt (M(unknown)/32.0) = 2

(M(unknown)/32.0) = 4

M(unknown) = 1.28*10^2 g/mol

Answer: E